4 Ω 3Ω 8V 5i, 1V 1 1Ω 2 Ω write all nodal equations for the circuit below
Answer #1

40) nog va 20 Ub SAVC 12 2131232 () 8) 4 ป (2) +sor 0 -1+44 + 3 ) + (1-ம்)=0 பிட் - 34 -ம் = 1 0 ட் - டி-=-(1)- 3 (உ-ம்) - 8+ 2 (i) +(i-i) = 0 3in +3* -புப் - - put in உம் in-Datei Page: put in @3iz+313 =41 12-1) = 0 1512+15:13 - 412+4*3= 0 1112 +1913 = 0 from xn & Q x7 -7112=14313=55 Aziz + 13313=